40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 00:07:54 2015, rana, Update, LSC, Lock Loss plot 1107673198.png1107673198.png
    Reply  Wed Feb 11 02:42:05 2015, Jenne, Update, LSC, PRC error signal RF spectra REFL33_25.pdfREFL33_15.pdf
       Reply  Wed Feb 11 03:41:12 2015, Koji, Update, LSC, PRC error signal RF spectra 
          Reply  Wed Feb 11 17:31:11 2015, ericq, Update, LSC, RFPD spectra REFL.pngAS.pngPOP.png2015-02-PDSpectra.zip
             Reply  Wed Feb 11 18:07:42 2015, ericq, Update, LSC, RFPD spectra 
                Reply  Thu Feb 12 01:00:18 2015, rana, Update, LSC, RFPD spectra 
Message ID: 10999     Entry time: Wed Feb 11 02:42:05 2015     In reply to: 10998     Reply to this: 11000
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRC error signal RF spectra 

Since we're having trouble keeping the PRC locked as we reduce the CARM offset, and we saw that the POP22 power is significantly lower in the 25% MICH offset case than without a MICH offset, Rana suggested having a look at the RF spectra of the REFL33 photodiode, to see what's going on. 

The Agilent is hooked up to the RF monitor on the REFL33 demod board.  The REFL33 PD has a notch at 11MHz and another at 55MHz, and a peak at 33MHz. 

We took a set of spectra with MICH at 25% offset, and another set with MICH at 15% offset.  Each of these sets has 4 traces, each at a different CARM offset.  Out at high CARM offset, the arm power vs. CARM offset is pretty much independent of MICH offset, so the CARM offsets are roughly the same between the 2 MICH offset plots. 

What we see is that for MICH offset of 25%, the REFL33 signal is getting smaller with smaller CARM offset!!  This means, as Rana mentioned earlier this evening, that there's no way we can hold the PRC locked if we reduce the CARM offset any more. 

However, for the MICH offset 15% case, the REFL 33 signal is getting bigger, which indicates that we should be able to hold the PRC.  We are still losing PRC lock, but perhaps it's back to mundane things like actuator saturation, etc. 

The moral of the story is that the 3f locking seems to not be as good with large MICH offsets.  We need a quick Mist simulation to reproduce the plots below, to make sure this all jives with what we expect from simulation.

For the plots, the blue trace has the true frequency, and each successive trace is offset in frequency by a factor of 1MHz from the last, just so that it's easier to see the individual peak heights.

Here is the plot with MICH at 25% offset:

And here is the plot with MICH at 15% offset:

Note that the analyzer was in "spectrum" mode, so the peak heights are the true rms values.  These spectra are from the monitor point, which is 1/10th the value that is actually used.  So, these peak heights (mVrms level) times 10 is what we're sending into the mixer.  These are pretty reasonable levels, and it's likely that we aren't saturating things in the PD head with these levels. 

The peaks at 100MHz, 130MHz and 170MHz that do not change height with CARM offset or MICH offset, we assume are some electronics noise, and not a true optical signal.

Also, a note to Q, the new netgpib scripts didn't write data in a format that was back-compatible with the old netgpib stuff, so Rana reverted a bunch of things in that directory back to the most recent version that was working with his plotting scripts.  sorry.

 

Attachment 1: REFL33_25.pdf  61 kB  | Hide | Hide all
REFL33_25.pdf
Attachment 2: REFL33_15.pdf  61 kB  | Hide | Hide all
REFL33_15.pdf
ELOG V3.1.3-